OpenStudy (anonymous):

Please help! :) Factor the trinomial. x^2 - 6x + 8

OpenStudy (anonymous):

Okay, to factor this we need to find 2 numbers that multiply to give positive 8, yet add to give negative 6. Well we know two negatives= a positive, so it will be 2 negative numbers. And here -2*-4=8 and -2-4=-6 \((x-2)(x-4)\)

OpenStudy (precal):

another way to look at it is you are looking for 8 1,2,4,8 the addition sign in front of 8 tells you to add the factors up 1+8=9 2+4=6 this is why we pick 2 and 4 now sign choices are -(-) or +(+) since they produce a positive result your middle number is -6x that is why we pick (-)(-) (x-2)(x-4)

OpenStudy (anonymous):

thank you both :)

OpenStudy (precal):

anytime

OpenStudy (anonymous):

That is a cool way to think about it, @precal my advanced algebra teacher taught me the way I did it years ago and I have used it since

OpenStudy (precal):

use whatever works for you....just wanted to give another way to look at it